# URLOST: Unsupervised Representation Learning without Stationarity or Topology

### v1

@@ Abstract
Unsupervised representation learning has seen tremendous progress but is constrained by its reliance on data modality-specific stationarity and topology, a limitation not found in biological intelligence systems. For instance, human vision processes visual signals derived from irregular and non-stationary sampling lattices yet accurately perceives the geometry of the world. We introduce a novel framework that learns from high-dimensional data lacking stationarity and topology. Our model combines a learnable self-organizing layer, density adjusted spectral clustering, and masked autoencoders. We evaluate its effectiveness on simulated biological vision data, neural recordings from the primary visual cortex, and gene expression datasets. Compared to state-of-the-art unsupervised learning methods like SimCLR and MAE, our model excels at learning meaningful representations across diverse modalities without depending on stationarity or topology. It also outperforms other methods not dependent on these factors, setting a new benchmark in the field. This work represents a step toward unsupervised learning methods that can generalize across diverse high-dimensional data modalities.
@@ 1 Introduction
Unsupervised representation learning, also known as self-supervised representation learning (SSL), aims to develop models that autonomously detect patterns in data and make these patterns readily apparent through a specific representation. There has been tremendous progress over the past few years in the unsupervised representation learning community. Popular methods like contrastive learning and masked autoencoders [68; 6; 24; 70] work relatively well on typical modalities such as images, videos, audio, time series, and point clouds. However, these methods make implicit assumptions about the data domain’s topology and stationarity. Given an image, topology refers to the neighboring pixels of each pixel, or more generally, the grid structure in images, the temporal structure in time series and sequences, or the 3D structure in molecules and point clouds. Stationarity refers to the property that the low-level statistics of the signal remain consistent across its domain. For instance, pixels and patches in images exhibit similar low-level statistics (mean, variance, co-variance) regardless of their locations within the domain. The success of state-of-the-art self-supervised representation learning relies on knowing the prior topology and stationarity of the modalities. For example, joint-embedding SSL employs random-resized cropping augmentation [6], and masked auto-encoding [25] utilizes masked-image-patch augmentation. What if we possess high-dimensional signals without knowledge of their domain topology or stationarity? Can we still craft a high-quality representation? This is not only the situation that biological vision systems have to deal with but also a practical setting for many scientific data analysis problems. In this work, we introduce unsupervised representation learning without stationarity or topology (URLOST) and take a step in this direction.
@@
As we mentioned earlier, typical modalities possess topology and stationarity prior information that can be utilized by unsupervised representation learning. Taking images as an example, digital cameras employ a consistent sensor grid that spans the entire visual field. However, biological visual systems have to deal with signals with less domain regularity. For instance, unlike camera sensors which have a uniform grid, the cones and rods in the retina distribute unevenly and non-uniformly. This results in a non-stationary raw signal. Retinal ganglion cells connect to more photoreceptors in the fovea than in the periphery. The correlation of the visual signal between two different locations in the retina depends not only on the displacement between these locations but also on their absolute positions. Yet, biological visual systems can establish precise retinotopy from the retina to neurons based on spontaneous retinal activities and external stimuli [67; 35; 18] and leverage retinotopic input to build unsupervised representation. This implies that we can potentially build unsupervised representation without relying on prior stationarity of the raw signal or topology of the input domain.
@@
Figure 1: From left to right: the unsupervised representation learning through joint embedding and masked auto-encoding; the biological vision system that perceives via unstructured sensor and understands signal without stationarity or topology [48]; and many more such diverse high dimensional signal in natural science that our method supports while most existing unsupervised methods don’t. Data figures are borrowed from [48; 44; 69].
@@
In this work, we aim to build unsupervised representations for general high-dimensional vectors. Taking images as an example again, let’s assume we receive a set of images whose pixels are shuffled in the same order. How can we build representations in an unsupervised fashion without knowledge of the shuffling order? If possible, can we use such a method to build unsupervised representations for general high-dimensional data? Inspired by [53], we use low-level statistics and spectral clustering to form clusters of the pixels, which recovers a coarse topology of the input domain. These clusters are analogous to image patches except that they are slightly irregularly shaped and different in size. We mask a proportion of these “patches” and utilize a Vision Transformer [15] to predict the masked “patches” based on the remaining unmasked ones. This “learning to predict masked tokens” approach is proposed in masked autoencoders (MAE) [25] and has demonstrated effectiveness on typical modalities. Initially, we test the proposed method on the synthesized biological visual dataset, derived from CIFAR-10 [32] using a foveated retinal sampling mechanism [8]. Then we generalize this method to two high-dimensional vector datasets: a primary visual cortex neural response decoding dataset [57] and the TCGA miRNA-based cancer classification dataset [62; 66]. Across all these benchmarks, our proposed method outperforms existing SSL techniques, establishing its effectiveness in building unsupervised representations for signals lacking explicit stationarity or topology. Given the emergence of new modalities in deep learning from natural sciences [59; 23; 49; 34], such as chemistry, biology, and neuroscience, our method offers a promising approach in the effort to build unsupervised representations for high-dimensional data.
@@ 2.1 Motivation and Overall framework
Our objective is to build robust, unsupervised representations for high-dimensional signals that lack explicit topology and stationarity. These learned representations are intended to enhance performance in downstream tasks, such as classification. To achieve this, we begin by using low-level statistics and clustering to approximate the signal’s topology. The clusters derived from the signal serve as input to a masked autoencoder. As depicted in Figure 1, the masked autoencoder randomly masks out patches in an image and trains a Transformer-based autoencoder unsupervisedly to reconstruct the original image. After the unsupervised training, the autoencoder’s latent state yields high-quality representations. In our approach, signal clusters are input to the masked autoencoder.
Notably, the clusters differ from image patches in several key aspects due to the differences in the input signal: they are unaligned, exhibit varied sizes and shapes, and their clustering nodes are not confined to fixed 2D locations like pixels in image patches. To cope with these differences, we introduce a self-organizing layer responsible for aligning these clusters through learnable transformations. The parameters of this layer are jointly optimized with those of the masked autoencoder. Our method is termed URLOST, an acronym for Unsupervised Representation Learning withOut Stationarity or Topology. Figure 2 provides an overview of the framework. URLOST consists of three core components: density adjusted spectral clustering, self-organizing layer, and masked autoencoder. The functionalities of these components are detailed in the following subsections.
@@ 2.2 Density Adjusted Spectral clustering
Representation learning for high-dimensional signals without explicit topology is challenging. We propose to define a metric to measure inter-dimensional relationships. This metric effectively approximates a topology for the signal. Similar to [53], where they use the absolute correlation values as the metric for pixels, we employ discrete mutual information (refer to Appendix A.1) as the metric. Let affinity matrix $A_{ij}$ denote the mutual information between dimension $i$ and $j$, which approximates the manifold $\mathcal{M}$ that the signal lives on. We can define the discretized Laplacian operator based on $A$ and use the eigenvector of the Laplacian operator to perform spectral clustering, which segments the manifold. The detailed definition and the algorithm are left in Appendix A.1. Finding the eigenvector of the Laplacian operator is a discretized approximation of the following optimization problem in function space:
$\min_{||f||_{L^{2}(\mathcal{M})}}\int_{\mathcal{M}}||p(x)^{\frac{1}{2}}\nabla f(x)||^{2}$ (1)
where $f(x):\mathcal{M}\to[0,1]$ is the normalized signal defined on $\mathcal{M}$ and $p(x)$ is the density function. The integral is taken over standard measure on $\mathcal{M}$. Since spectral clustering heavily relies on the solution of equation 1, the definition of the density function $p(x)$ affects the quality of the resulting clusters. Standard approaches often assume that nodes are uniformly distributed on the manifold, thereby treating $p(x)$ as a constant and excluding it from the optimization process. However, this assumption does not hold in our case involving non-stationary signals. To cope with non-stationary signals, our work introduces a variable density function $p(x)$ for each signal, making it a pivotal component in building good representations for the signal. This component is referred to as Density Adjusted Spectral Clustering. Empirical evidence supporting this design is provided through visualization and ablation studies in the experimental section.
@@ 2.3 Self-organizing layer
Transforming a high-dimensional signal into a sequence of clusters using the above method is not enough because it does not capture the internal structure within individual clusters. To effectively perform unsupervised learning on these clusters, it is essential to align them in some manner. Directly solving the exact alignment problem with low-level statistics of the signal is challenging. Thus, we propose a self-organizing layer with learnable parameters. Specifically, let vector $x^{(i)}$ denote the $i$th cluster. Each cluster $x^{(i)}$ is passed through a differentiable function $g(\cdot,w^{(i)})$ with parameter $w^{(i)}$, resulting in a sequence $z_{0}$:
@@ 2.4 Masked autoencoder
After the self-organizing layer, $z_{0}$ is passed to a Transformer-based masked autoencoder (MAE) with an unsupervised learning objective. Masked autoencoder (MAE) consists of an encoder and a decoder which both consist of stacked Transformer blocks introduced in [64]. The objective function is introduced in [25]: masking random image patches in an image and training an autoencoder to reconstruct them, as illustrated in Figure 1. In our case, randomly selected clusters in $z_{0}$ are masked out, and the autoencoder is trained to reconstruct these masked clusters. After training, the encoder’s output is treated as the learned representation of the input signal for downstream tasks. The masked prediction loss is computed as the mean square error (MSE) between the values of the masked clusters and their corresponding predictions.
@@ 3 Result
Since our method is inspired by the biological vision system, we first validate its ability on a synthetic biological vision dataset created from CIFAR-10. Then we evaluate the generalizability of URLOST on two high-dimensional natural datasets collected from diverse domains. Detailed information about each dataset and the corresponding experiments is presented in the following subsections. Across all tasks, URLOST consistently outperforms other strong unsupervised representation learning methods.
@@ 3.1 Synthetic biological vision dataset
As discussed in the introduction, the biological visual signal serves as an ideal dataset to validate the capability of URLOST. In contrast to digital images captured by a fixed array of sensors, the biological visual signal is acquired through irregularly positioned ganglion cells, inherently lacking explicit topology and stationarity. However, it is hard to collect real-world biological vision signals with high precision. Therefore, we employ a retinal sampling technique to modify the classic CIFAR-10 dataset and simulate imaging from the biological vision signal. The synthetic dataset is referred to as Foveated CIFAR-10. To make a comprehensive comparison, we also conduct experiments on the original CIFAR-10, and a Permuted CIFAR-10 dataset obtained by randomly permuting the image.
@@
Foveated CIFAR-10. Much like photosensors installed in a camera, retina ganglion cells within the primate biological visual system sample from visual stimuli and project images. However, unlike photosensors that have uniform receptive fields and adhere to a consistent sampling pattern, retinal ganglion cells at different locations of the retina vary in their receptive field size: smaller in the center (fovea) but larger in the peripheral of the retina. This distinctive retina sampling pattern results in foveated imaging [63]. It gives primates the ability to have both a high-resolution vision and a broad overall receptive field while consequently making visual signals sampled by the retina lack stationarity. The evidence is that responses of two ganglion cells separated by the same displacement are highly correlated in the retina but less correlated in the peripheral. To mimic the foveated imaging with CIFAR-10, we adopt the retina sampling mechanism from [8]. Specifically, each retina ganglion cell is simplified and modeled using a Gaussian kernel. The response of each cell is determined by the dot product between pixel values and the Gaussian kernel. Figure 3 illustrates the sampling kernel locations. Applying this sampling grid and permuting the resulting pixels produces the foveated CIFAR-10. In the natural retina, retinal ganglion cell density decreases linearly with eccentricity, which makes fovea much denser than the peripheral, compared to the simulated lattice in Figure 3. However, considering the low resolution of the CIFAR-10 dataset, we reduce the simulated fovea’s density to prevent redundant sampling.
@@
Experiments. We compare URLOST on both of the synthetic vision datasets as well as the original CIFAR-10 with popular unsupervised representation learning methods SimCLR [6] and MAE [25]. All the models conducted unsupervised learning followed by linear probing for classification accuracy. The evaluations are reported in Table 1. SimCLR excels on CIFAR-10 but struggles badly with both synthetic datasets due to its inability to handle data without stationarity and topology. MAE gets close to SimCLR on CIFAR-10 with a $4\times 4$ patch size. However, the patch size no longer makes sense when data has no topology. So we additionally tested MAE masking pixels instead of image patches. It maintains the same performance on Permuted CIFAR-10 as on CIFAR-10, though poorly, invariant to the removal of topology as it should be. But It still drops greatly to $48.5\%$ on the Foveated CIFAR-10 when stationarity is also removed. In contrast, only URLOST is able to maintain consistently strong performances when there is no topology or stationarity, achieving $86.4\%$ on Permuted CIFAR-10 and $85.4\%$ on Foveated CIFAR-10 when the baselines completely fail.
@@
Table 1: Evaluation on computer vision and synthetic biological vision dataset. ViT (Patch) stands for the Vision Transformer backbone with image patches as inputs. ViT (Pixel) means pixels are treated as input units. ViT (Clusters) means clusters are treated as inputs instead of patches. The number of clusters is set to 64 for both Permuted CIFAR-10 and Foveated CIFAR-10 dataset.
@@
Table 2: Evaluation on V1 response decoding and TCGA pan-cancer classification tasks. “Raw” indicates preprocessed (standardized and normalized) raw signals. Best $\beta$ values are used for $\beta$-VAE. For URLOST MAE, cluster sizes are 200 (V1) and 32 (TCGA).

### v2

@@ Abstract
Unsupervised representation learning has seen tremendous progress. However, it is constrained by its reliance on domain specific stationarity and topology, a limitation not found in biological intelligence systems. For instance, unlike computer vision, human vision can process visual signals sampled from highly irregular and non-stationary sensors. We introduce a novel framework that learns from high-dimensional data without prior knowledge of stationarity and topology. Our model, abbreviated as URLOST, combines a learnable self-organizing layer, spectral clustering, and a masked autoencoder (MAE). We evaluate its effectiveness on three diverse data modalities including simulated biological vision data, neural recordings from the primary visual cortex, and gene expressions. Compared to state-of-the-art unsupervised learning methods like SimCLR and MAE, our model excels at learning meaningful representations across diverse modalities without knowing their stationarity or topology. It also outperforms other methods that are not dependent on these factors, setting a new benchmark in the field. We position this work as a step toward unsupervised learning methods capable of generalizing across diverse high-dimensional data modalities. Code is available at this repository.
@@ 1 Introduction
Unsupervised representation learning aims to develop models that autonomously detect patterns in data and make these patterns readily apparent through specific representations. Over the past few years, there has been tremendous progress in the unsupervised representation learning community, especially self-supervised representation learning (SSL) method. Popular methods SSL methods like contrastive learning and masked autoencoding [87; 12; 33; 90] work well on typical data modalities such as images, videos, time series, and point clouds. However, these methods make implicit assumptions about the data domain’s topology and stationarity. Topology refers to the low-dimensional structure arisen from physical measurements, such as the pixel grids in images, the temporal structures in time series and text sequences, or the 3D structures in molecules and point clouds [7]. Stationarity refers to the characteristic that the statistical properties of the signal are invariant across its domain  [6]. For instance, the statistics of pixels and patches in images are invariant to their spatial locations. A vase is still a vase no matter placed at the corner or the center of the image.
The success of state-of-the-art self-supervised representation learning (SSL) methods largely depends on these two crucial assumptions. For example, in computer vision, popular SSL techniques, such as Masked Autoencoders [34] and joint-embedding methods like SimCLR [12] require the construction of image patches. Both approaches typically rely on convolutional neural networks (CNNs) or Vision Transformer (ViT) backbones [22]. These backbones consist of shared-weight convolutional filters or linear layers, which inherently exploit the stationarity and regular topology present in natural images. The geometric deep learning community has made significant efforts to extend machine learning to domains beyond those with regular topology. However, graph neural network (GNN)-based methods empirically struggle to scale with self-supervised learning objectives and large datasets [69; 10; 9]. Methods that do scale well with large data still assume a minimal level of stationarity and regular topology [32].
@@
What if we have high-dimensional signals without prior knowledge of their domain topology or stationarity? Can we still craft a high-quality representation? This is not only the situation that biological intelligence systems have to deal with but also a practical setting for many scientific data analysis problems. Taking images as an example, computer vision system takes in signal from digital cameras. Biological visual systems, on the other hand, have to deal with signals with less domain regularity. Unlike the uniform grid in camera sensors, the cones and rods in the retina are distributed unevenly and non-uniformly. Yet, biological visual systems can establish a precise retinotopy map from the retina to neurons in visual cortex based on spontaneous, locally-propagated retinal activities and external stimuli [86; 47; 25] and leverage retinotopic input to build unsupervised representations. This motivates us to build unsupervised representations without relying on the prior stationarity of the raw signal or the topology of the input domain. The ability to build unsupervised representations without relying on topology and stationarity has huge advantages. For example, it allows humans to utilize irregular sensors for both high resolution and broad coverage, which is much more efficient than a regular camera sensor for dynamic environments. Developing such a model also allows us to create a powerful AI system that computes any high-dimensional signal.
@@
Figure 1: From left to right: the unsupervised representation learning through joint embedding and masked auto-encoding; the biological vision system that perceives via unstructured sensor and understands signal without stationarity or topology [61]; and many more such diverse high dimensional signal in natural science [57; 89] that our method supports while most existing unsupervised methods don’t.
@@
In this work, we aim to build unsupervised representations for general high-dimensional data and introduce unsupervised representation learning without stationarity or topology (URLOST). Taking images as an example again, let’s assume we receive a set of images whose pixels are shuffled in the same order. In this case, the original definition of topology of images is destroyed, i.e. each pixel should no longer the neighbor of the pixels that’s physically next to it. How can we build representations in an unsupervised fashion without knowledge of the shuffling order? If possible, can we use such a method to build unsupervised representations for general high-dimensional data? Inspired by Roux et al. [67], we use low-level statistics and spectral clustering to form clusters of the pixels, which recovers a coarse topology of the input domain. These clusters are analogous to image patches except that they are slightly irregularly shaped and different in size. We mask a proportion of these “patches” and utilize a Vision Transformer [22] to predict the masked “patches” based on the remaining unmasked ones. This “learning to predict masked tokens” approach is proposed in masked autoencoders (MAE) [34] and has demonstrated effectiveness on typical modalities. Firstly, we test the proposed method on the synthesized biological visual dataset, derived from the CIFAR-10 [44] using a foveated retinal sampling mechanism [14]. Then we generalize this method to two high-dimensional vector datasets: a primary visual cortex neural response decoding dataset [73] and the TCGA miRNA-based cancer classification dataset [79; 84]. Across all these benchmarks, our proposed method outperforms existing SSL methods, demonstrating its effectiveness in building unsupervised representations for signals lacking explicit stationarity or topology. Given the emergence of new modalities in deep learning from natural sciences [75; 31; 62; 46; 85; 17; 2], such as chemistry, biology, and neuroscience, our method offers a promising approach in the effort to build unsupervised representations for high-dimensional data.
@@ 2.1 Motivation and Overall Framework
Our objective is to build robust unsupervised representations for high-dimensional signals without prior information on explicit topology and stationarity. The learned representations are intended to enhance performance in downstream tasks such as classification. We begin by using low-level statistics and clustering to approximate the topology of the signal domain. The clusters then serve as input to a masked autoencoder. As depicted in Figure 1, the masked autoencoder randomly masks out patches in an image and trains a Transformer-based autoencoder unsupervisedly to reconstruct the original image. After unsupervised training, the latent state of the autoencoder yields high-quality representations. In our approach, the clusters derived from the raw signal are input to the masked autoencoder.
Taking images as an example data modality, clusters formed in images are bags of pixels, which differs from image patches in several key aspects. First, they are unaligned, exhibit varied sizes and shapes. Second, each pixel in a cluster is not confined to fixed 2D locations like pixels in image patches. To cope with these differences, we introduce a self-organizing layer responsible for aligning these clusters through learnable transformations. The parameters of this layer are jointly optimized with those of the masked autoencoder. Our method is termed URLOST, an acronym for Unsupervised Representation Learning withOut Stationarity or Topology. Figure 2 provides an overview of the framework. URLOST consists of three core components: density adjusted spectral clustering, self-organizing layer, and masked autoencoder. The functionalities of these components are detailed in the following subsections.
@@ 2.2 Density Adjusted Spectral Clustering
Given a dataset $S\in\mathbb{R}^{n\times m}$, where each row denote a high dimensional data point. Let $S_{i}\in\mathbb{R}^{n\times 1}$ be $i$th column of $S$, which represents the $i$th dimension of the signal. Since we want to process dimensions that share similar information together, we use spectral clustering to group these dimensions. Let $A_{ij}=I(S_{i};S_{j})$ be the mutual information between dimension $i$ and $j$. Consider each dimension as a node of a graph and $A$ is the the affinity matrix of a graph. $L=D-A$ is the graph Laplacian, where D is the diagonal matrix whose $(i,i)$-element is the sum of $A$’s $i$-th row. We can formulate the spectral embedding problem as finding $Y$ such that $\min_{YY^{T}=I}tr(YLY^{T})$. A clustering algorithm is then applied to the embedding $Y$ as explained in appendix A.2. The size and shape of the clusters strongly affect the unsupervised learning performance. To adjust the size and shape, we apply a density adjustment matrix $P$ to adjust $L$ in the spectral embedding objective. The optimization problem becomes the following:
$\min_{YY^{T}=I}tr(YP^{1/2}LP^{1/2}Y^{T})$ (1)
where $P=diag(p(i))$, $p(i)$ is the unnormalized density function defined on each node $i$. We set $p(i)=q(i)^{\alpha}n(i)^{-\beta}$ , where $n(i)=\sum_{j\in\text{Top}_{K}(A_{ji})}A_{ji}$ and $q(i)$ is the prior density which depends on specific dataset and is defined in the experiment section. $\alpha$, $\beta$ and $K$ are hyper-parameters. Setting $\alpha=0$ and $K=m$ will recover the normalized graph Laplacian. In appendix A.1, we provide a detailed interpretation and motivation of density adjustment. We further verify its effectiveness with ablation study in section 4.3 and appendix A.4
@@ 2.3 Self-organizing Layer
Transforming a high-dimensional signal into a sequence of clusters using the above method is not enough because it does not capture the internal structure within individual clusters. As an intuitive example, given an image, we divide it into a set of image patches of the same size. If we apply different permutations to these image patches, their inner product will no longer reflect their similarity properly. Clusters we obtained from section 2.2 are analogous to image patches, but elements in each cluster have arbitrary ordering. Thus, the inner product between two clusters is also arbitrary due to the ordering mismatch. In Transformers, since self-attention depends on the inner products between different “clusters,” we need to align these clusters in a space, where their inner products reflect their similarity. To align these clusters, we propose a self-organizing layer with learnable parameters. Specifically, let vector $x^{(i)}$ denote the $i$th cluster. Each cluster $x^{(i)}$ is passed through a differentiable function $g(\cdot,w^{(i)})$ with parameter $w^{(i)}$, resulting in a sequence $z_{0}$:
@@ 2.4 Masked Autoencoder
After the self-organizing layer, $z_{0}$ is passed to a Transformer-based masked autoencoder (MAE) with an unsupervised learning objective. Masked autoencoder (MAE) consists of an encoder and a decoder which both consist of stacked Transformer blocks introduced in Vaswani et al. [82]. The objective function is introduced in He et al. [34]: masking random image patches in an image and training an autoencoder to reconstruct them, as illustrated in Figure 1. In our case, randomly selected clusters in $z_{0}$ are masked out, and the autoencoder is trained to reconstruct these masked clusters. After training, the encoder’s output is treated as the learned representation of the input signal for downstream tasks. The masked prediction loss is computed as the mean square error (MSE) between the values of the masked clusters and their corresponding predictions.
@@ 3 Result
Since the biological vision system inspires our method, we first validate its ability on a synthetic biological vision dataset created from CIFAR-10. Then we evaluate the generalizability of URLOST on two high-dimensional natural datasets collected from diverse domains. Detailed information about each dataset and the corresponding experiments is presented in the following subsections. Across all tasks, URLOST consistently outperforms other strong unsupervised representation learning methods. For all the experiments in this work, we use linear layer to parametrize the self-organizing layer, i.e. $g(x,W^{(i)})=W^{(i)}x$. Additionally, we provide training hyperparameters and experiments for effect of hyperparameters in Appendix A.6.
@@ 3.1 Synthetic Biological Vision Dataset
As discussed in the introduction, the biological visual signal serves as an ideal data modality to validate the capability of URLOST. In contrast to digital images captured by a fixed array of sensors, the biological visual signal is acquired through irregularly positioned ganglion cells, inherently lacking explicit topology and stationarity. However, it is hard to collect real-world biological vision signals with high precision and labels to evaluate our algorithm. Therefore, we employ a retinal sampling technique to modify the classic CIFAR-10 dataset and simulate imaging from the biological vision signal. The synthetic dataset is referred to as Foveated CIFAR-10. To make a comprehensive comparison, we also conduct experiments on the original CIFAR-10, and a Permuted CIFAR-10 dataset obtained by randomly permuting the image.
@@
Foveated CIFAR-10. Unlike cameras that use uniform photosensors with consistent sampling patterns, the human retina features ganglion cells whose receptive field sizes vary—smaller in the central fovea and larger in the periphery. This results in foveated imaging [80], enabling primates to have both high-resolution central vision and a wide overall receptive field. However, this distinctive sampling pattern causes visual signals in the retina to lack stationarity; the statistical properties differ between the center and periphery, with ganglion cell responses being highly correlated in the fovea but less correlated in the periphery. To mimic the foveated imaging with CIFAR-10, we adopt the retina sampling mechanism from Cheung et al. [14]. Like shown in Figure 3, each dot represents a retinal ganglion cell, which together form a non-stationary sampling lattice with irregular topology. Details on the implementation of foveation sampling is provided in Appendix A.3.
@@
Experiments. We compare URLOST on both of the synthetic vision datasets as well as the original CIFAR-10 with popular unsupervised representation learning methods SimCLR [12] and MAE [34]. We also compared the ViT backbone used in URLOST to other backbones such as convolutional neural network (CNN). Moreover, since CNN works better with stationary signals, we further compared our methods to SimCLR with a graph neural network (GNN) backbone, which doesn’t rely on the stationarity assumption of the data. For the GNN baseline, we use Vision GNN (ViG), which is state of the art graph neural network on images proposed in [32]. After the model is trained without supervision, we use linear probing to evaluate it. This is achieved by training a linear classifier on top of the pre-trained model with the given labels. The evaluations are reported in Table 1. SimCLR excels on CIFAR-10 but struggles with both synthetic datasets due to its inability to handle data without stationarity and topology. MAE gets close to SimCLR on CIFAR-10 with a $4\times 4$ patch size. However, the patch size no longer makes sense when data has no topology. So we additionally tested MAE masking pixels instead of image patches. The performance of pixel level MAE is invariant to the removal of topology. The performance on Permuted CIFAR-10 is the same as the performance on CIFAR-10, though poorly. But it still drops greatly to $48.5\%$ on the Foveated CIFAR-10 when stationarity is also removed. In contrast, only URLOST is able to maintain consistently strong performances when there is no topology or stationarity, achieving $86.4\%$ on Permuted CIFAR-10 and $85.4\%$ on Foveated CIFAR-10 when the baselines completely fail.
@@
Table 1: Evaluation on computer vision and synthetic biological vision dataset. ViT (Patch) stands for the Vision Transformer backbone with image patches as inputs. ViT (Pixel) means pixels are treated as input units. ViT (Clusters) means clusters are treated as inputs instead of patches. The number of clusters is set to 64 for both Permuted CIFAR-10 and Foveated CIFAR-10 dataset. Eval Acc stands for linear probing evaluation accuracy.
@@
SimCLR ViG 42.8 %
Table 2: Evaluation on V1 response decoding and TCGA pan-cancer classification tasks. “Raw” indicates preprocessed (standardized and normalized) raw signals. Best $\beta$ values are used for $\beta$-VAE. For URLOST MAE, cluster sizes are 200 (V1) and 32 (TCGA). We pick 15 seeds randomly to repeat the training and evaluation for each method. For the MAE baseline, we randomly group dimensions instead of using spectral clusterings. We report the 95% confidence interval for all methods.